# Photoacoustic Iterative Optimization Algorithm with Shape Prior Regularization

### v1

@@ Abstract
Photoacoustic imaging (PAI) suffers from inherent limitations that can degrade the quality of reconstructed results, such as noise, artifacts, and incomplete data acquisition caused by sparse sampling or partial array detection. In this study, we proposed a new optimization method for both two-dimensional (2D) and three-dimensional (3D) PAI reconstruction results, called regularized iteration method with shape prior. The shape prior is a probability matrix derived from the reconstruction results of multiple sets of random partial array signals in a computational imaging system using any reconstruction algorithm, such as Delay-and-Sum (DAS) and Back-Projection (BP). In the probability matrix, high-probability locations indicate high consistency among multiple reconstruction results at those positions, suggesting a high likelihood of representing the true imaging results. In contrast, low-probability locations indicate higher randomness, leaning more towards noise or artifacts. As a shape prior, this probability matrix guides the iteration and regularization of the entire array signal reconstruction results using the original reconstruction algorithm (the same algorithm for processing random partial array signals). The method takes advantage of the property that the similarity of the object to be imitated is higher than that of noise or artifact in the results reconstructed by multiple sets of random partial array signals of the entire imaging system. The probability matrix is taken as a prerequisite for improving the original reconstruction results, and the optimizer is used to further iterate the imaging results to remove noise and artifacts and improve the imaging fidelity. Especially in the case involving sparse view which brings more artifacts, the effect is remarkable. Simulation and real experiments have both demonstrated the superiority of this method.
@@ 1 Introduction
      In this study, we proposed a new optimization method for photoacoustic imaging reconstruction results. The key idea of this method is that the similarity of the imaged object is higher than the noise and artifacts in results of the reconstruction of multiple random partial array signals in the entire imaging array using any reconstruction algorithm, such as DAS and BP, and using these results a probability matrix is obtained through calculation methods such as difference. The different positions of the matrix reflect the probability that the position is the real object in the imaging space. As a shape prior[17], this probability matrix guides the iteration and regularization of the entire array signal reconstruction results using the original reconstruction algorithm (the same reconstruction algorithm for processing random partial array signals). This method is not affected by the array distribution, imaging region, etc., and can be used in both 2D and 3D imaging in photoacoustic computational imaging, and can be widely applied to other computational imaging systems, including ultrasound, ct, etc., with strong generalization and robustness. By leveraging the shape prior based on possibility matrix and iteratively refining reconstruction results, this method achieves superior performance compared to the result before the optimization. Both simulation and experimental results demonstrate the effectiveness of the proposed method in reducing noise and artifacts under sparse-view conditions while preserving critical structural information in PAI reconstructions. For the convenience of calling this algorithm, we call it RI-SPPM, RI means that it is a regularized iteration process, SPPM means that the shape prior is based on a probability matrix.
@@ 2.1 Construction of Probability Matrix
This section begins by introducing the causes of noise and artifacts in photoacoustic imaging, laying the foundation for the rationale of the proposed method. For the construction of the probability matrix, we randomly select a portion of the signal from all array elements for reconstruction. Among the obtained reconstruction results, one result is chosen as the reference baseline, and the others are compared with the reference result by performing pixel-wise difference calculations. A probability score is then defined based on the proportion of these differences, resulting in a probability matrix that matches the dimensionality of the reconstruction results. In this matrix, high-probability regions indicate a higher likelihood of representing the true imaged object, while noise and artifacts, due to their randomness, tend to correspond to regions with lower probabilities. Finally, we describe how this shape prior knowledge is utilized in a specific regularized iterative reconstruction process, including the selection of the loss function and the optimizer.
@@ 2.1.2 Reconstruction of Random Selected Element Signals
For a computational imaging system with array number $N$ and sampling number $M$, the obtained signal domain data dimension is $N*M$. We randomly select $x$ element signals for reconstruction ($x<N$ and the reconstruction algorithm chooses the algorithm you want to optimize the results), here we have A selections(A = $c_{N}^{x}$), and we only need A random k results (k<A), the k results are $R_{1},R_{2},...R_{k}$.(The choices of $x$ and $k$ are described in detail in subsequent paragraphs)
@@ 2.1.3 Acquisition of Probability Matrix
If we take $R_{1}$ as the base, then the probability matrix of the $i$ $th$ result with that base can be defined as $P_{i}$.
$P_{i}=1-\frac{\left|R_{i}-R_{1}\right|}{\left|R_{1}\right|}$ (1)
This matrix reflects the similarity relationships of individual voxel/pixel. Other calculation methods or mixed approaches can also be used. For example, the similarity of a 3×3×3/3×3 region can be used to represent the similarity at the center position, or a weighted calculation can be performed by combining regional similarity and individual position similarity. It should be noted in particular that we do not process some data in $R$ where the position itself is 0, because it represents that the position has no noise and no imaged object, and the probability value has little significance, and it does not need to be over-processed in the regular iteration process. At the same time, the numerical significance of the position less than 0 in $P_{i}$ indicates that there is a great difference in the data of this position. However, in order to have actual probability significance, we take a value of 0 to indicate that the probability of this place being an imaged object is extremely low.
So our final probability matrix is:
$P=\frac{P_{2}+...+P_{k}}{k}$ (2)
@@ 2.2 Regularized Iteration Optimization
In this section, we will use the probability matrix $P$ obtained in the previous section as a shape prior to design a suitable loss function for regular iterative optimization of the original reconstruction results of the entire array signal.
@@ 2.2.1.1 Shape Prior
It is not difficult to find that $P$ is a probability matrix distributed in $0-1$. But in order to have a more intuitive physical meaning for the subsequent regular iteration process, the region with a value close to $1$ (representing the high confidence part of the shape prior, the optimization goal is to maintain or expand the value of these regions).Regions with a sum value close to 00 (indicating regions of low confidence or no interest for which the optimization goal is to impose forced suppression) are more sensitive, and we normalize $P$ $0-1$. The result after normalization is $P_{N}$.
${P_{N}=\frac{P-min(P)}{max(P)-min(P)}}$ (3)
2D Shape Prior
For two-dimensional reconstruction results, the probability matrix in RI-SPPM is a two-dimensional matrix with the same dimensions as the reconstruction images, which reflects the difference of pixels at the corresponding positions of the reconstruction results of multiple random array signals, that is, the pixels at a high probability position represent a higher probability of being the real object to be imitated with strong consistency. The low probability position means that the pixels in this position are more inclined to random noise or artifacts.
3D Shape Prior
For the three-dimensional reconstruction results, the probability matrix in RI-SPPM is a three-dimensional matrix with the same dimensions as the reconstruction results, which reflects the difference of voxels in the corresponding positions of the reconstruction results of multiple random element signals, that is, the voxels in the position with high probability represent a higher probability of being the real object to be imaged with strong consistency. A low probability position means that the voxel at that position is more biased towards noise or artifacts with strong randomness.
The feasibility of using $P_{N}$ as a shape prior lies in the fact that it directly provides the probability distribution information of each voxel or pixel region, which can be used to indicate the model’s tendency to strengthen certain regions (high probability regions) in the reconstruction results and its inhibition demand for other regions (low probability regions). Such prior information can effectively guide the optimization process, making the result more in line with the expected structural characteristics, while reducing the impact of noise or artifacts. The rational constraint of shape information is realized mathematically, and the optimization stability of shape consistency are improved. It has clear feasibility and physical significance.
@@ 2.2.1.2 Loss Function
${gradient_{dc}=\frac{\partial{Loss}_{dc}}{\partial R_{op}}=2*(R_{N}-R_{op})}$ (5)
@@
${gradient_{rg}=\frac{\partial{Loss}_{rg}}{\partial R_{op}}=2*(1-P_{N})*R_{op}}$ (7)
@@ 2.2.2.2 Regularized Iteration Process
10:               $gradient_{data\_consistency}\leftarrow 2*(R_{N}-R_{op})$
@@
12:               $gradient_{prior\_reg}\leftarrow 2*(1-P_{N})*R_{op}$
@@ 3.1 Simulation Data Results
We first used simulated hand vessels as the target tissue to demonstrate the high performance of the RI-SPPM algorithm. The simulated data is based on the 3D PA imaging results of human peripheral hand vessels from another work[22], which was used as the ground truth. At a grid point spacing of 0.2 mm, the total grid size of the imaging area is 128*320*256 (25.6 mm ×64.0 mm ×51.2 mm). In order to simulate 3D PA imaging, the radius of the hemispherical system is 4cm, and a gap with a height of 1cm is opened at the top as the laser light inlet, and the remaining positions are evenly distributed with 1024 array elements(Figure 3 is imaging setup). Then we obtain PA signals collected by these 1024 arrays through forward propagation of the K-Wave toolbox[23].
@@
The above imaging system is selected for simulation to generate both real and sparse photoacoustic signals for evaluating the performance of the RI-SPPM algorithm. The simulation is based on 3D photoacoustic imaging data of real hand blood vessels as the "ground truth", and adopts a high-precision grid with a resolution of 0.2mm (128×320×256) to capture the complex structure of blood vessels. The system is designed as a hemispherical array with a radius of 4 cm and a laser inlet with a height of 1 cm at the top, and 1024 sensors are uniformly distributed on the array to simulate a typical photoacoustic imaging device. This simulation method can approximate the actual conditions, provide controllable evaluation criteria, and provide a reliable basis for verifying the performance of RI-SPPM algorithm in PAI.
3.1.2 Comparison of Results before and after Optimization
In this example, we use the 3D PA result reconstructed by the UBP algorithm as the result optimized by the RI-SPPM algorithm. The acoustic source distributions are presented in Figure 4(a), while the reconstruction outcomes are depicted in Figure 4(b-c). For each set of images, the upper and lower sub-images correspond to the top-view maximum amplitude projection (MAP) and the front-view MAP of the 3D reconstruction, respectively. In Figure 4(b), the reconstruction results obtained via the UBP algorithm display significant artifacts on the right side of the index finger. These arise because the laser illumination originates from the right side of the index finger, leading to a lack of information from that particular angle and, consequently, severe artifacts. When compared to the ground truth, a substantial discrepancy is evident. As shown in Figure 4(c), the optimization performed by the RI-SPPM algorithm significantly reduces the artifacts in the UBP results, bringing them closer to the ground truth. The artifacts caused by incomplete angular coverage are largely corrected, and additional spurious artifacts in other regions are also mitigated, such as those observed in the middle of the finger’s top-view MAP and in the upper and lower areas of the finger’s front-view MAP. Furthermore, in the 3D view (Figure 4(e-f)), the improvements in the PA reconstruction results after applying the RI-SPPM algorithm are clearly observable.
@@ 3.1.2 Comparison of Results before and after Optimization
(b) Finger’s top-view maximum amplitude projection, front-view maximum amplitude projection of the UBP results.
(c) Finger’s top-view maximum amplitude projection, front-view maximum amplitude projection of the RI-SPPM results. (d) Reconstruction result of UBP algorithm in 3D-view. (e) Reconstruction result after optimization of RI-SPPM algorithm in 3D-view.
Subsequently, we performed a comprehensive mean squared error (MSE) analysis on the reconstruction results before and after optimization (all three datasets are normalized). The MSE between the reconstruction using the UBP algorithm and the ground truth is $0.00270$, whereas the MSE after applying the RI-SPPM optimization is reduced to $0.00078$. To further demonstrate the effectiveness of the RI-SPPM optimization, and considering that the overall MSE metric may be influenced by the presence of artifacts, we selected a region rich in valid information for detailed analysis and verification. Specifically, we extracted data along a line defined in dimensions a and b. The voxel intensity values along this line for the three datasets are presented in Figure 5. From Figure 5(a), it can be observed that, in regions lacking finger data, the noise and artifact levels in the RI-SPPM-optimized results are significantly closer to zero, while the noise and artifact levels in the UBP-only reconstruction are considerably higher. In Figure 5(b), it is evident that, in regions containing finger information, the intensity curve of the RI-SPPM-optimized result closely aligns with the ground truth, demonstrating superior consistency compared to the UBP-only reconstruction.
(a) No finger information
@@
Figure 5: (a) shows the Line 64, :,230 intensity distribution without finger information. (b) shows line 64, :,190 intensity distribution with finger information
@@
3.1.3 Parameters Selection and Computing Environment
@@
3.2 Vivo Animal Experiment Results

### v2

@@ Abstract
Photoacoustic imaging (PAI) suffers from inherent limitations that can degrade the quality of reconstructed results, such as noise, artifacts and incomplete data acquisition caused by sparse sampling or partial array detection. In this study, we proposed a new optimization method for both two-dimensional (2D) and three-dimensional (3D) PAI reconstruction results, called the regularized iteration method with shape prior. The shape prior is a probability matrix derived from the reconstruction results of multiple sets of random partial array signals in a computational imaging system using any reconstruction algorithm, such as Delay-and-Sum (DAS) and Back-Projection (BP). In the probability matrix, high-probability locations indicate high consistency among multiple reconstruction results at those positions, suggesting a high likelihood of representing the true imaging results. In contrast, low-probability locations indicate higher randomness, leaning more towards noise or artifacts. As a shape prior, this probability matrix guides the iteration and regularization of the entire array signal reconstruction results using the original reconstruction algorithm (the same algorithm for processing random partial array signals). The method takes advantage of the property that the similarity of the object to be imitated is higher than that of noise or artifact in the results reconstructed by multiple sets of random partial array signals of the entire imaging system. The probability matrix is taken as a prerequisite for improving the original reconstruction results, and the optimizer is used to further iterate the imaging results to remove noise and artifacts and improve the imaging fidelity. Especially in the case involving sparse view which brings more artifacts, the effect is remarkable. Simulation and real experiments have both demonstrated the superiority of this method.
@@ 1 Introduction
      In this study, we proposed a new optimization method for photoacoustic imaging reconstruction results. The key idea of this method is that the similarity of the imaged object is higher than the noise and artifacts in results of the reconstruction of multiple random partial array signals in the entire imaging array using any reconstruction algorithm, such as DAS and BP, and using these results a probability matrix is obtained through calculation methods. The different positions of the matrix reflect the probability that the position is the real object in the imaging space. As a shape prior[17], this probability matrix guides the iteration and regularization of the entire array signal reconstruction results using the original reconstruction algorithm (the same reconstruction algorithm for processing random partial array signals). This method is not affected by the array distribution, imaging region, etc., and can be used in both 2D and 3D imaging in photoacoustic computational imaging, and can be widely applied to other computational imaging systems, including ultrasound, ct, etc., with strong generalization and robustness. By leveraging the shape prior based on possibility matrix and iteratively refining reconstruction results, this method achieves superior performance compared to the result before the optimization. Both simulation and experimental results demonstrate the effectiveness of the proposed method in reducing noise and artifacts under sparse-view conditions while preserving critical structural information in PAI reconstructions. For the convenience of calling this algorithm, we call it RI-SPPM, RI means that it is a regularized iteration process, SPPM means that the shape prior is based on a probability matrix.
@@ 2.1 Construction of Probability Matrix
This section begins by introducing the causes of noise and artifacts in photoacoustic imaging, laying the foundation for the rationale of the proposed method. For the construction of the probability matrix, we randomly select a portion of the signal from all array elements for reconstruction. A probability matrix consistent with the dimensions of the reconstruction results is generated using mathematical operations on the obtained reconstruction results. In this matrix, high-probability regions indicate a higher likelihood of representing the true imaged object, while noise and artifacts, due to their randomness, tend to correspond to regions with lower probabilities. Finally, we describe how this shape prior knowledge is utilized in a specific regularized iterative reconstruction process, including the selection of the loss function and the optimizer.
@@ 2.1.2 Reconstruction of Random Selected Element Signals
For a computational imaging system with array number $N$ and sampling number $M$, the obtained signal domain data dimension is $N*M$. We randomly select $x$ element signals for reconstruction ($x<N$), here we have A selections(A = $c_{N}^{x}$), and we only need A random k results ($k<A$), the $k$ results are $R_{1},R_{2},...R_{k}$. (The choices of $x$ and $k$ are described in detail in subsequent paragraphs.)
@@ 2.1.3 Acquisition of Probability Matrix
In three dimensions, the probability matrix is defined as $P(x,y,z)$.
$P(x,y,z)=\int_{v_{\min}}^{v_{\max}}D(x,y,z)\cdot\frac{v-v_{\min}}{v_{\max}-v_{\min}}\cdot\|D(x,y,z)-v>0\|_{0}\mathrm{d}v$ (1)
where, $v_{min}=min(D(x,y,z))$, $v_{max}=max(D(x,y,z))$
$D(x,y,z)=\frac{(\frac{1}{k}(R_{1}+R_{2}+...+R_{k}))^{2}}{\frac{1}{k}(R_{1}^{2}+R_{2}^{2}+...R_{k}^{2})}=\frac{(R_{1}+R_{2}+...+R_{k})^{2}}{{k}\times(R_{1}^{2}+R_{2}^{2}+...R_{k}^{2})}$ (2)
In order to have a more intuitive physical meaning for the subsequent regular iteration process, we normalize $P$ $0-1$. The result after normalization is $P_{N}$.
${P_{N}=\frac{P-min(P)}{max(P)-min(P)}}$ (3)
The function $P(x,y,z)$ is introduced as a probabilistic representation of spatial data distributions in three-dimensional space. The function integrates the local property $D(x,y,z)$, which quantifies the concentration or uniformity of measured attributes $R_{1},R_{2},...,R_{k}$ at a given point $(x,y,z)$, defined as the ratio of the squared arithmetic mean to the mean of squared values. This makes $D(x,y,z)$ a measure of localized data consistency, where higher values indicate greater uniformity. To account for variations across the defined region, $P(x,y,z)$ is computed by integrating over the range $[v_{m}in,v_{m}ax]$, where these bounds represent the global extremal values of $D(x,y,z)$. The integration involves a weighting factor $\frac{v-v_{\min}}{v_{\max}-v_{\min}}\cdot\|D(x,y,z)-v\|_{0}$, which characterizes the probabilistic "importance" of this point within the entire three-dimensional region through weighting and integration range.
@@ 2.2 Regularized Iteration Optimization
In this section, we will use the probability matrix $P$ obtained in the previous section as a shape prior and design a suitable loss function for regular iterative optimization of the original reconstruction results of the entire array signal.
@@ 2.2.1.1 Shape Prior
The feasibility of using $P_{N}$ as a shape prior lies in the fact that it directly provides the probability distribution information of each voxel or pixel region, which can be used to indicate the model’s tendency to strengthen certain regions (high probability regions) in the reconstruction results and its inhibition demand for other regions (low probability regions). This prior information can effectively guide the optimization process, making the result more in line with the expected structural characteristics, while reducing the impact of noise or artifacts. The rational constraint of shape information is realized mathematically, and the optimization stability of shape consistency is improved. It has clear feasibility and physical significance.
@@ 2.2.1.2 Loss Function
${gradient_{dc}=\frac{\partial{Loss}_{dc}}{\partial R_{op}}=2\times(R_{N}-R_{op})}$ (5)
@@
${gradient_{rg}=\frac{\partial{Loss}_{rg}}{\partial R_{op}}=2\times(1-P_{N})\times R_{op}}$ (7)
@@ 2.2.2.2 Regularized Iteration Process
10:               $gradient_{data\_consistency}\leftarrow 2\times(R_{N}-R_{op})$
@@
12:               $gradient_{prior\_reg}\leftarrow 2\times(1-P_{N})\times R_{op}$
@@ 3.1 Simulation Data Results
We first used simulated hand vessels as the target tissue to demonstrate the high performance of the RI-SPPM algorithm. The simulated data is based on the 3D PA imaging results of human peripheral hand vessels from another work[22], which was used as the ground truth. At a grid point spacing of 0.2 mm, the total grid size of the imaging area is 128*320*256 (25.6 mm ×64.0 mm ×51.2 mm). In order to simulate 3D PA imaging, the radius of the hemispherical system is 4cm, and a gap with a height of 1cm is opened at the top as the laser light inlet, and the remaining positions are evenly distributed with 1024 array elements(Figure 3 is imaging setup). Then we obtained PA signals collected by these 1024 arrays through forward propagation of the K-Wave toolbox[23].
@@
The above imaging system was selected for simulation to generate sparse photoacoustic signals in order to evaluate the performance of the RI-SPPM algorithm. The simulation is based on 3D photoacoustic imaging data of real hand vasculature as the "ground truth," using a high-resolution grid with a resolution of 0.2 mm (128×320×256) to capture the intricate structures of the vasculature. The system is designed as a hemispherical array with a radius of 4 cm, featuring a laser entry point at the top with a height of 1 cm to create a missing-view angle. The array is uniformly distributed with 1024 sensors to simulate a typical photoacoustic imaging device. This simulation approach closely approximates real-world scenarios, provides a controlled evaluation standard, and offers a reliable basis for validating the performance of the RI-SPPM algorithm in photoacoustic imaging.
3.1.1.1 Comparison of Results before and after Optimization
In this example, we use the 3D PA result reconstructed by the UBP algorithm as the result optimized by the RI-SPPM algorithm. The distributions of the acoustic source are presented in Figure 4(a), while the reconstruction results are depicted in Figure 4(b-c). For each set of images, the upper and lower subimages correspond to the top-view maximum amplitude projection (MAP) and the front-view MAP of the 3D reconstruction, respectively. In Figure 4(b), the reconstruction results obtained via the UBP algorithm display significant artifacts on the right side of the index finger. These arise because the laser illumination originates from the right side of the index finger, leading to a lack of information from that particular angle and, consequently, severe artifacts. When compared to the ground truth, a substantial discrepancy is evident. As shown in Figure 4(c), the optimization performed by the RI-SPPM algorithm significantly reduces the artifacts in the UBP results, bringing them closer to the ground truth. The artifacts caused by incomplete angular coverage are largely corrected, and additional spurious artifacts in other regions are also mitigated, such as those observed in the middle of the finger’s top-view MAP and in the upper and lower areas of the finger’s front-view MAP. Furthermore, in the 3D view (Figure 4(e-f)), the improvements in the PA reconstruction results after applying the RI-SPPM algorithm are clearly observable.
@@ 3.1.1.1 Comparison of Results before and after Optimization
(b) Finger’s top-view maximum amplitude projection, front-view maximum amplitude projection of the UBP result.
(c) Finger’s top-view maximum amplitude projection, front-view maximum amplitude projection of the RI-SPPM result. (d) Reconstruction result of UBP algorithm in 3D-view. (e) Reconstruction result after optimization of RI-SPPM algorithm in 3D-view.
Subsequently, we performed a comprehensive mean squared error (MSE) analysis on the reconstruction results before and after optimization (all three datasets are normalized). The MSE between the reconstruction using the UBP algorithm and the ground truth is $0.3398$, whereas the MSE after applying the RI-SPPM optimization is reduced to $0.0040$. To further demonstrate the effectiveness of the RI-SPPM optimization, and considering that the overall MSE metric may be influenced by the presence of artifacts, we selected a region rich in valid information for detailed analysis and verification. Specifically, we extracted data along two lines. The voxel intensity values along the lines for the three datasets are presented in Figure 5. In Figure 5(b), it is evident that, in regions containing finger information, the intensity curve of the RI-SPPM-optimized result aligns closely with the ground truth, demonstrating superior consistency compared to the UBP-only reconstruction. The correlations before and after using RI-SPPM are $0.65$ and $0.92$. In Figure 5 (c), it can be seen that in regions lacking finger data, the noise and artifact levels in the RI-SPPM-optimized results are significantly closer to zero, while the noise and artifact levels in the UBP-only reconstruction are considerably higher. The correlations before and after using RI-SPPM are $0.00$ and $0.87$.
(a) The location of the two lines used for data analysis
@@
(c) No finger information
Figure 5: (a) Red line with the finger information and blue line without the finger information. (b) shows the intensity distribution in red line. (c) shows the intensity distribution in blue line.
@@
3.1.1.2 Parameters Selection and Computing Environment
@@
3.1.2 3D PA Image Reconstruction under 256 Spherical System
We then used a Gaussian sphere and three cylinders as target to demonstrate the high performance of the RI-SPPM algorithm (Figure 6(a)). At a grid point spacing of 0.1 mm, the total grid size of the imaging area is 256*256*256 (25.6 mm ×25.6 mm ×25.6 mm). In order to simulate 3D PA imaging, the radius of the spherical system is 6cm. Then we obtained PA signals collected by these 256 arrays through forward propagation of the K-Wave toolbox.
3.1.2.1 Comparison of Results before and after Optimization
From a three-dimensional perspective (Figure 6(b,c)), the artifacts significantly decrease after optimization by the RI-SPPM algorithm. The Gaussian sphere and tubular structures are displayed more clearly. Through calculations, the MSE between the UBP reconstructed result and the ground truth is $0.0126$, while the MSE of the result optimized by the RI-SPPM algorithm is $0.000146$. In conclusion, in the simulation experiments, the reconstruction results obtained using the UBP algorithm and further optimized by the RI-SPPM algorithm demonstrated excellent performance, particularly in artifact suppression, noise reduction, and the preservation and representation of true structural information.
(a) Ground truth
(b) UBP reconstruction
(c) RI-SPPM optimization
Figure 6: Comparison of 3D photoacoustic reconstruction results under 256 spherical system. (a) Ground truth. (b) Reconstruction result of UBP algorithm in 3D-view. (c) Reconstruction result after optimization of RI-SPPM algorithm in 3D-view.
To better show the contrast, in Figure 7 we use a jet map to display the maximum amplitude projection in the direction of the red arrow in Figure 6(a).
(a) Ground truth
(b) UBP reconstruction
(c) RI-SPPM optimization
Figure 7: Comparison of 3D photoacoustic reconstruction results under 256 spherical system. (a) The maximum amplitude projection of ground truth. (b) The maximum amplitude projection of the UBP result. (c) The maximum amplitude projection of the RI-SPPM result.
3.1.2.2 Parameters Selection and Computing Environment
Parameters Value
$N$ 256
$M$ 4096
$x$ 25
$k$ 50
$\lambda_{con}$ 0.10
$\lambda_{reg}$ 0.90
$lr$ 0.001
$num\_iters$ 500
Table 2: 3D PA image reconstruction under 256 spherical system
In this experiment, we used a graphics card type GeForce RTX 4090 and the CPU type is AMD® Epyc 9354 32-core processor x 128. In UBP algorithm reconstruction, we use Taichi for GPU acceleration. For the reconstruction with PA signal size of 256*4096 and region size of 256*256*256, the time consuming is about 0.8s. So, the entire RI-SPPM optimization process took less than a minute and 30 seconds. Of course, we have skills in the choice of parameters, and the choice of x is generally about one-tenth of the total number of elements. We also found that when the initial value of lr is about 0.001, the iterative result can be quickly optimized. The number of iterations is much lower than num_iters.
3.2 In Vivo Human Experiment Results